A particle of charge 3.00nC is located at point (0,0) in a certain coordinate system. Assuming all lengths are in meters, calculate the potential at the following points: A (3,3) and B (0,2)


1
Expert's answer
2021-02-07T19:22:31-0500

The potential from a point charge at a distance "r" from it is given as follows:


"\varphi = k\dfrac{q}{r}"


where "q = 3\times 10^{-9}C" is the charge, and "k = 9\times 10^9N\cdot m^2/C^2" is the constant.

In Cartesian coordinates the distance from the origin to some point is:


"r = \sqrt{x^2 + y^2}"

where "x,y" are coordinates of this point. Thus, case of point "A (3,3)" have:


"\varphi_A = k\dfrac{q}{\sqrt{x^2 + y^2}} = \dfrac{9\times 10^9\cdot 3\times 10^{-9}}{\sqrt{3^2 + 3^2}} =4.5\sqrt{2} \approx 6.36V\\"

Similarly, for point "B(0,2)" obtain:


"\varphi_B = k\dfrac{q}{\sqrt{x^2 + y^2}} = \dfrac{9\times 10^9\cdot 3\times 10^{-9}}{\sqrt{0^2 + 2^2}} =13.5V\\"

Answer. "\varphi_A = 4.5\sqrt{2} \approx 6.36V, \varphi_B = 13.5V" .